# Scene-aware Human Motion Forecasting via Mutual Distance Prediction

### v3

@@ Abstract
In this paper, we tackle the problem of scene-aware 3D human motion forecasting. A key challenge of this task is to predict future human motions that are consistent with the scene by modeling the human-scene interactions. While recent works have demonstrated that explicit constraints on human-scene interactions can prevent the occurrence of ghost motion, they only provide constraints on partial human motion e.g., the global motion of the human or a few joints contacting the scene, leaving the rest of unconstrained. To address this limitation, we propose to model the human-scene interaction with the mutual distance between the human body and the scene. Such mutual distances constrain both the local and global human motion, resulting in a whole-body motion constrained prediction. In particular, mutual distance constraints consist of two components, the signed distance of each vertex on the human mesh to the scene surface and the distance of basis scene points to the human mesh. We further introduce a global scene representation learned from a signed distance function (SDF) volume to ensure coherence between the global scene representation and the explicit constraint from the mutual distance. We develop a pipeline with two sequential steps: predicting the future mutual distances first, followed by forecasting future human motion. During training, we explicitly encourage consistency between predicted poses and mutual distances. Extensive evaluations on the existing synthetic and real datasets demonstrate that our approach consistently outperforms the state-of-the-art methods.
@@ 1 Introduction
Human motion forecasting has a wide range of applications in human robot interaction [15], autonomous driving [23], virtual/augmented reality (AR/VR) [29] and animation [31]. Despite being an active research field for decades [2, 28, 30, 33], most previous methods [17, 1, 22, 32, 8, 16, 21, 3] focus solely on human motion and overlook the fact that humans interact with the 3D environment. By contrast, this paper tackles the problem of scene-aware human motion forecasting.
Recent works in human motion prediction [4, 5, 40, 20] and scene-aware human synthesis [34, 37, 11] have begun to model the human-scene interaction. Some works do so implicitly with the scene encoding as an additional input to their model e.g., 2D scene image [4], 3D scene point cloud [34] or a specific object [5]. However, without explicit constraints, those methods tend to predict motions with artifacts e.g., “ghost motion”. Alternatively, other works [37, 40, 20, 27] start to explore the explicit representation of human-scene interaction. However, none of their representations are able to constrain the whole-body motion both locally and globally. For example, the proximity between human and scene used in [37] mainly constrains the global position of the human, while the human gaze employed in [40] only specifies the potential target location of the human. Although the contact map proposed in [20, 27] achieves detailed constraints on both the local pose and global position of certain human joints, their contact map can only constrain joints that contact the scene, neglecting the others, such as elbow joints, which often remain unsupported.
@@
To address this, we leverage a per-basis point distance, which calculates the minimum distance between each point in a predefined set of basis points and the human surface. Our per-basis point distance is similar to the BPS body encoding in PLACE [37] with differences in the choices of basis points. In particular, their basis points are partially scene-dependent making it not suitable for motion prediction, while our basis points are purely scene-independent and sampled with a fixed distance relative to the origin, which can better capture the motion patterns.
@@
Our contributions can be summarized as follows. i)We propose to explicitly represent the human-scene interaction as the mutual distance to provide constraints for the whole-body motion. ii) We explore different global scene representations and find the SDF volume to be more efficient than commonly used point cloud representation. Our experiments on both synthetic and real datasets demonstrate that our approach outperforms the state-of-the-art scene-aware human motion forecasting methods by a large margin on existing benchmark datasets.
@@ 2 Related Work
Human motion prediction. Researchers have been studying 3D human motion modeling [2, 28, 30, 33] for decades. While periodic and simple non-periodic motions like walking and golf swings can be handled by traditional methods such as those using Hidden Markov Models [2] or the Gaussian process latent variable model [33], more complex motions are typically modeled with deep neural networks [17, 1, 22, 32, 8, 16, 21, 3]. Those methods however only predict the local poses ignoring the global human motion and do not take the 3D scenes where the motion is performed into account.
@@
More recent works [4, 5, 20, 40, 27] have started to explore the scene context in the task of motion prediction. In particular, Cao et al. [4] propose to represent the scene with a 2D image. The feature extracted from such an image is fed into the prediction model and lets the model learn the human-scene interaction implicitly. However, such 2D scene representation cannot handle occlusions and it does not ensure a valid human-scene interaction. In [5], a semantic graph is introduced to encode the interaction between the human and the object of interest. Although it is efficient to model a human interacting with one object, it is not suitable for human-scene interaction as the human may interact with multiple objects when he moves around. In [40], Zheng et al. employ eye gaze as additional information to guide the motion forecasting process. Although eye gaze provides valuable hints about the goal location of the human motion, it constrains neither the local poses nor the human-scene interaction. Mao et al. [20] introduce a per-joint contact map that can explicitly encode the relation between scene surface points and human joints. Luca et al. [27] propose a three-stage pipeline in a coarse-to-fine manner and utilizing the time information based on [20]. However, for forecasting, their contact map only constrains joints that contact the scene and disregard the rest. By contrast, our introduced mutual distance is able to constrain the whole-body motion.
@@
Scene-aware human synthesis. One of the main challenges in this task is to ensure a valid human-scene interaction. As such, existing works have explored various strategies to model such interaction [37, 39, 11, 35, 34, 9]. Similar to [4], the interaction is modeled implicitly in [39, 35, 34] via either Generative Adversarial Networks (GANs) [34], or Variational Autoencoders [39, 34]. Besides this, other works [37, 11, 9] adopt the explicit representation of human-scene interaction. In particular, it was proposed in POSA [11] to learn a human-centric representation of the human-scene interaction which assigns scene semantic labels, such as floor, and chair, to each human surface vertex. Those labels indicate the probability of such vertex contacting what kind of object in the scene. Although it performs well in the task of putting a human in a scene, it does not provide much information about the local pose of the human, thus is ill-suited for human motion forecasting. In SAMP [9], Hassan et al. propose to only model the interaction with the object at the goal location. Given the final root translation and orientation, the task is then treated as a path planning problem. Similar to [5], such a strategy cannot constrain the interactions between the human and the scene when the human moves in the scene. The closest to our work is PLACE [37] where the human-scene interaction is represented by the minimum distance between a set of 3D scene points also known as Basis Point Sets (BSP) [26] to the human body surface. However, such representation alone only provides a prior on the global position of the human. For example, one can move its arm that is far from the scene surface while the scene to human proximity remains the same. By contrast, we combine the per-basis point distance with the per-vertex signed distance resulting in a detailed constraint on the whole-body motion both locally and globally.
@@ 3 Approach
Let us first formulate the task of scene-aware 3D human motion forecasting. Following previous work [20], a sequence of $T$ past human poses is represented as $\mathbf{X}=[\mathbf{x}_{1},\mathbf{x}_{2},\cdots,\mathbf{x}_{T}]\in\mathbb{R}^{M\times T}$ [^1] Unless otherwise specified, in this paper, we use column vector convention., where a human pose at time $t$ is represented as a vector $\mathbf{x}_{t}\in\mathbb{R}^{M}$ of length $M$. Such vector could be a parametric representation of human mesh such as SMPL-X [25] pose. To represent the 3D scene, we adopt a Signed Distance Volume $\mathbf{S}\in\mathbb{R}^{N\times N\times N}$, each entry of which stores the distance of the voxel to the nearest scene surface and its sign indicates whether it is inside or outside the scene. The representation learned from this volume is more consistent with our mutual distance representation than those learned from point cloud [20]. Moreover, such Signed Distance Volume allows us to query the signed distance of any given position easily via interpolation, which facilitates the fast calculation of the loss defined on the SDFs of the predicted pose sequence during training (see Eq. 14).
Given the historical motion $\mathbf{X}$ and the 3D scene $\mathbf{S}$, our goal is then to forecast $U$ future human poses $\mathbf{Y}=[\mathbf{x}_{T+1},\mathbf{x}_{T+2},\cdots,\mathbf{x}_{T+U}]\in\mathbb{R}^{M\times U}$. To this end, we introduce a novel pipeline that first predicts the mutual distance representation i.e., the signed minimum distance from each human body vertex to the scene surface (the per-vertex signed distance) and the minimum distance from every basis point to the human surface (the per-basis point distance) at each time frame. Conditioned on such mutual distance, our model then forecasts the future poses. An overview of our pipeline is shown in Fig. 2. Below, we present our mutual distance representation and our pipeline to predict these distances and future motions.
Figure 2: Network architecture. Given the 3D scene $\mathbf{S}$ represented as a signed distance volume and the past motion $\mathbf{X}$ shown in grey meshes, our approach first predicts the future per-vertex signed distance (top row), and the future per-basis point distance (bottom row) from historical distances. The predicted distances are then fed into a RNN-based network to predict the future motion shown in orange meshes (middle row).
@@ 3.1 Mutual Distance Representation
where $\mathbf{d}_{t}=[d_{1t},d_{2t},\cdots,d_{Kt}]^{T}\in\mathbb{R}^{K}$ is the per-vertex signed distance at time step $t$, $\mathcal{S}$ is the set of the points occupied by the scene and $\partial\mathcal{S}$ is the surface points of such scene. The per-vertex signed distance provides information about the whole-body pose. However, the same human mesh at different global positions may result in the same per-vertex signed distance. To resolve such ambiguity, we further leverage the per-basis point distance which is described below.
@@ 3.2 Mutual Distance Prediction
where $\mathcal{G}_{b}$ is also a GCN. $\hat{\mathbf{K}}$ is the DCT coefficients of the padded past per-basis point distance sequence. $\hat{\mathbf{B}}\in\mathbb{R}^{P\times(T+U)}$ is the predicted per-basis point distance.
@@
where $\hat{d}_{kt}$ is the predicted signed distance for the $k$-th vertex at time $t$ and $\hat{b}_{pt}$ is the predicted distance for the $p$-th basis point at time $t$, $d_{kt}$ and $b_{pt}$ define the ground truth.
@@ 3.3 Motion Forecasting Network
Here we adopt a stage-wise training strategy where we initially train the mutual distance prediction network and motion forecasting network separately. While our network is differentiable and can be trained from scratch in an end-to-end fashion, we observed that at the early stage of the training process, the learning process can easily converge to a local minimum due to the inaccurate prediction of the mutual distance which serves as the input for the motion forecasting network. We thus propose to perform stage-wise training first. Note that, for the stage-wise training process, the motion forecasting network is provided with the ground truth mutual distance as input. Subsequently, we combine both networks and fine-tune them in an end-to-end manner with both losses defined in Eq. 10 and 16.
@@ 4.1 Datasets
GTA-IM. The GTA Indoor Motion dataset [4], building upon GTA game engine, is a synthetic human-scene interaction dataset. It comprises approximately 1 million RGB-D frames featuring 3D human poses for 50 distinct characters across 7 diverse scenes. To reconstruct the mesh, we fuse the depth maps of each scene and subsequently generate the SDF volume. We follow ContactAware [20] to use the same training-test split and 21 out of the 98 human joints. The videos run at 30 FPS. Both our model and baseline models are trained to observe the past 30 frames to predict the future 60 frames following [20].
@@
PROX. Proximal Relationships with Object eXclusion (PROX) [10] is a real dataset with 20 subjects interacting with the scenes. The dataset provides human SMPL-X parameters [25] fitted from RGB-D frames. Since the original motion sequences are jittery, we use the refined version provided by [20]. We adopt the same data split and the 22 body joints of SMPL-X model as in [20]. The frame-rate is 30 Hz. Our model and baseline models take the past 30 frames as input to predict the future 60 frames following [20].
@@
HUMANISE. The HUMANISE dataset [36] is a large-scale synthetic dataset that captures human-scene interactions. It is mainly built upon two primary datasets: the AMASS dataset [19] and the ScanNet dataset [6]. HUMANISE combines these datasets by incorporating different global motions to meet contact and collision constraints. Since the original split includes duplicated human motions in the training and test set which is not suitable for motion prediction, we create a new split to ensure that motions in the test set are never observed during training. Moreover, we ensure that the scenes paired with the motions in the test set are categorized into seen scenes, which include a set of scenes observed during the training, and unseen scenes consisting of scenes never observed during training. Our final split comprises approximately 20k sub-sequences for training and 6k for testing, which includes 5k sub-sequences from seen scenes and 1k from unseen scenes. The videos run at 30 frame-per-second (FPS). Both our model and baseline models are trained to observe the past 15 frames (0.5 sec.) and predict the future 30 frames (1 sec.) because the average length of motion sequences in the dataset is 2.2 sec. which is much shorter than 3 sec.
@@
GIMO. The Gaze-Informed Human Motion dataset (GIMO) [40] is a real human-scene interaction dataset. The dataset captures human motion using an IMU-based motion capture system, and the 3D scene is scanned using a smartphone equipped with lidar sensors. Same as [36], GIMO also adopts SMPL-X [25] representation. It captures 11 subjects performing various actions e.g., open window, lie in bed, in 19 scenes, resulting in 127 long motion sequences with 192k frames. We follow the official training and test split, resulting in around 6k training sequences and 1k testing sequences. Same as HUMANISE [36], the frame rate is 30Hz. Our model and baseline models are provided with 15 past frames to predict future 30 frames.

### v4

@@ Abstract
In this paper, we address the issue of scene-aware 3D human motion forecasting. A key challenge in this task is to predict future human motions that are coherent with the scene by modeling human-scene interactions.  While recent works have demonstrated that explicit constraints on human-scene interactions can prevent the occurrence of ghost motion, they only provide constraints on partial human motion e.g., the global motion of the human or a few joints contacting the scene, leaving the rest of unconstrained. To address this limitation, we propose to represent the human-scene interaction using the mutual distance between the human body and the scene. Such mutual distances constrain both the local and global human motion, resulting in a whole-body motion constrained prediction. In particular, mutual distance constraints consist of two components, the signed distance of each vertex on the human mesh to the scene surface and the distance of basis scene points to the human mesh. We further introduce a global scene representation learned from a signed distance function (SDF) volume to ensure coherence between the global scene representation and the explicit constraint from the mutual distance. We develop a pipeline with two sequential steps: predicting the future mutual distances first, followed by forecasting future human motion. We explicitly ensure consistency between predicted poses and mutual distances during training. Extensive testing on both synthetic and real datasets demonstrates that our method consistently surpasses the performance of current state-of-the-art techniques.
@@ 1 Introduction
Human motion forecasting has numerous applications, including autonomous driving [23], human-robot interaction [15], animation [31] and virtual/augmented reality [29]. Despite being an active research field for decades [2, 28, 30, 33], most previous methods [17, 1, 22, 32, 8, 16, 21, 3] focus solely on human motion and overlook the fact that humans interact with the 3D environment. In contrast, this paper addresses the problem of scene-aware human motion prediction.
Recent advancements in scene-aware human synthesis [34, 37, 11] and human motion prediction [4, 5, 40, 20] have begun to model the human-scene interaction. Some approaches achieve this implicitly by incorporating scene encoding as an additional input to their model e.g., 2D scene image [4], 3D scene point cloud [34] or a specific object [5]. However, without explicit constraints, those methods tend to predict motions with artifacts e.g., “ghost motion”. Alternatively, other works [37, 40, 20, 27] start to explore the explicit representation of human-scene interaction. However, none of their representations are able to constrain the whole-body motion both locally and globally. For example, the proximity between human and scene used in [37] mainly constrains the global position of the human, while the human gaze employed in [40] only specifies the potential target location of the human. Although the contact map proposed in [20, 27] achieves detailed constraints on both the local pose and global position of certain human joints, their contact map can only constrain joints that contact the scene, neglecting the others, such as elbow joints, which often remain unsupported.
@@
To address this, we leverage a per-basis point distance, which calculates the minimum distance between each point in a predefined set of basis points and the human surface. Our per-basis point distance is similar to BPS body encoding in PLACE [37] with differences in choices of basis points. In particular, their basis points are partially scene-dependent making it not suitable for motion prediction, while our basis points are purely scene-independent and sampled with a fixed distance relative to origin, which can better capture the motion patterns.
@@
Our contributions can be summarized as follows. i)We propose an explicit representation of human-scene interaction using mutual distance to provide constraints for whole-body motion. ii) We explore different global scene representations and find the SDF volume to be more efficient than commonly used point cloud representation. Our experiments demonstrate that our approach outperforms the state-of-the-art scene-aware human motion prediction methods by a large margin on both existing real and synthesis benchmark datasets.
@@ 2 Related Work
Human motion prediction. Researchers have been studying 3D human motion modeling [2, 28, 30, 33] for decades. While traditional methods, such as those employing Hidden Markov Models [2] or the Gaussian process latent variable model [33], are capable of handling periodic and simple non-periodic motions like walking and golf swings, more complex motions are typically modeled with deep neural networks [17, 1, 22, 32, 8, 16, 21, 3]. However, these methods focus solely on the prediction of local poses, neglecting the global human motion and 3D scenes in which the motion takes place.
@@
More recent works [4, 5, 20, 40, 27] have begun to incorporate scene context into the task of motion prediction. Specifically, Cao et al. [4] propose to represent the scene with a 2D image. The feature extracted from such an image is fed into the prediction model and lets the model learn the human-scene interaction implicitly. However, such 2D scene representation cannot handle occlusions and it does not ensure a valid human-scene interaction. In [5], a semantic graph is introduced to encode the interaction between the human and the object of interest. Although it is efficient to model a human interacting with one object, it is not suitable for human-scene interaction as the human may interact with multiple objects when he moves around. In [40], Zheng et al. employ eye gaze as additional information to guide the motion forecasting process. Although eye gaze provides valuable hints about the goal location of the human motion, it constrains neither the local poses nor the human-scene interaction. Mao et al. [20] introduce a contact map that can explicitly encode the relation between scene surface points and human joints. Scofano et al. [27] propose a three-stage pipeline in a coarse-to-fine manner and utilizing the time information based on [20]. However, for forecasting, their contact map only constrains joints that contact the scene and disregard the rest. By contrast, our introduced mutual distance is able to constrain the whole-body motion.
@@
Scene-aware human synthesis. One of the main challenges in this task is to ensure a valid human-scene interaction. As such, existing works have explored various strategies to model such interaction [37, 39, 11, 35, 34, 9]. Similar to [4], the interaction is modeled implicitly in [39, 35, 34] via either Generative Adversarial Networks (GANs) [34], or Variational Autoencoders [39, 34]. Besides this, other works [37, 11, 9] adopt explicit representation of human-scene interaction. In particular, it was proposed in POSA [11] to develop a human-centric representation of the human-scene interaction which assigns scene semantic labels, such as floor, and chair, to each human surface vertex. Those labels indicate the probability of such vertex contacting what kind of object in the scene. Although it performs well in the task of placing a human in a scene, it lacks sufficient detail about the local pose of the human, rendering it unsuitable for human motion forecasting. In SAMP [9], Hassan et al. propose to only model the interaction with the object at the goal location. Given the final root translation and orientation, the task is then treated as a path planning problem. Similar to [5], such a strategy cannot constrain the interactions between the human and the scene when the human moves in the scene. The closest to our work is PLACE [37] where the human-scene interaction is represented by the minimum distance between a set of 3D scene points also known as Basis Point Sets (BSP) [26] to the human body surface. However, such representation alone only provides a prior on the global position of the human. For example, one can move its arm that is far from the scene surface while the scene to human proximity remains the same. By contrast, we combine the per-basis point distance with the per-vertex signed distance resulting in a detailed constraint on the whole-body motion both locally and globally.
@@ 3 Approach
Let us begin by formulating the task of scene-aware 3D human motion forecasting. Building on previous work [20], we represent a sequence of $T$ past human poses as $\mathbf{X}=[\mathbf{x}_{1},\mathbf{x}_{2},\cdots,\mathbf{x}_{T}]\in\mathbb{R}^{M\times T}$ [^1] Unless otherwise specified, in this paper, we use column vector convention., where a human pose at time $t$ is represented as a vector $\mathbf{x}_{t}\in\mathbb{R}^{M}$ of length $M$. Such vector could be a parametric representation of human mesh such as SMPL-X [25] pose. To represent the 3D scene, we adopt a Signed Distance Volume $\mathbf{S}\in\mathbb{R}^{N\times N\times N}$, each entry of which stores the distance of the voxel to the nearest scene surface and its sign indicates whether it is inside or outside the scene. The representation learned from this volume is more consistent with our mutual distance representation than those learned from point cloud [20]. Moreover, such Signed Distance Volume allows us to query the signed distance of any given position easily via interpolation, which facilitates the fast calculation of the loss defined on the SDFs of the predicted pose sequence during training (see Eq. 14).
Our objective is predicting $U$ future human poses $\mathbf{Y}=[\mathbf{x}_{T+1},\mathbf{x}_{T+2},\cdots,\mathbf{x}_{T+U}]\in\mathbb{R}^{M\times U}$ with history motion $\mathbf{X}$ and 3D scene $\mathbf{S}$. We introduce an innovative pipeline that predicts mutual distance representation first i.e., the signed minimum distance from each human body vertex to the scene surface (the per-vertex signed distance) and the minimum distance from each basis point to the human surface (the per-basis point distance) at each time frame. Conditioned on such mutual distance, our model then forecasts future poses. Our pipeline is illustrated in Fig. 2. We detail our mutual distance and the pipeline for predicting these distances and future motions below.
Figure 2: Network architecture. Given the 3D scene $\mathbf{S}$ represented as a signed distance volume and the past motion $\mathbf{X}$ shown in grey meshes, our approach first predicts the future per-vertex signed distance $\hat{\mathbf{D}}$, and the future per-basis point distance $\hat{\mathbf{B}}$ from historical distance $\mathbf{D}$ and $\mathbf{B}$, respectively. The two predicted future distances are then fed into a RNN-based network to predict the future motion $\hat{\mathbf{Y}}$ shown in orange meshes.
@@ 3.1 Mutual Distance Representation
where $\mathbf{d}_{t}=[d_{1t},d_{2t},\cdots,d_{Kt}]^{T}\in\mathbb{R}^{K}$ is the per-vertex signed distance at time step $t$, $\mathcal{S}$ represents the volume inside the actual solid scene and $\partial\mathcal{S}$ is the surface points of such scene. The per-vertex signed distance provides information about the whole-body pose. However, the same human mesh at different global positions may result in the same per-vertex signed distance. To resolve such ambiguity, we further leverage the per-basis point distance which is described below.
@@ 3.2 Mutual Distance Prediction
where $\mathcal{G}_{b}$ is also a GCN. $\hat{\mathbf{K}}$ is the DCT coefficients of padded past per-basis point distance sequence. $\hat{\mathbf{B}}\in\mathbb{R}^{P\times(T+U)}$ is the predicted per-basis point distance.
@@
where $\hat{d}_{kt}$ is predicted signed distance for $k$-th vertex at time $t$ and $\hat{b}_{pt}$ is predicted distance for $p$-th basis point at time $t$, $d_{kt}$ and $b_{pt}$ define the ground truth.
@@ 3.3 Motion Forecasting Network
Here we adopt a stage-wise training strategy where we initially train the mutual distance prediction network and motion forecasting network separately. While our network is differentiable and can be trained from scratch in an end-to-end fashion, we observed that at early stage of the training process, the learning process can easily converge to a local minimum due to inaccurate prediction of mutual distance which serves as the input for the motion forecasting network. We thus propose to perform stage-wise training first. Note that, for stage-wise training process, the motion forecasting network is provided with the ground truth mutual distance as input. Subsequently, we combine both networks and fine-tune them in an end-to-end manner with both losses defined in Eq. 10 and 16.
@@ 4.1 Datasets
GTA-IM. The GTA Indoor Motion dataset [4], building upon GTA game engine, is a synthetic human-scene interaction dataset. It comprises approximately 1 million RGB-D frames featuring 3D human poses for 50 distinct characters across 7 diverse scenes. To reconstruct the mesh, we fuse the depth maps of each scene and subsequently generate the SDF volume. We follow ContactAware [20] to use the same training-test split and 21 joints. The videos are 30 FPS. Both our model and baseline models are trained to observe the past 30 frames to predict the future 60 frames following [20].
@@
PROX. Proximal Relationships with Object eXclusion (PROX) [10] is a real-world dataset. This dataset contains human SMPL-X parameters [25] fitted from RGB-D frames and 20 distinct characters interacting with various scenes. Since the original motion sequences are jittery, we use the refined version provided by [20]. We adopt the same data split and the 22 joints as in [20]. The frame rate is 30 Hz. Our model and baseline models input the past 30 frames to forecast the future 60 frames following [20].
@@
HUMANISE. The HUMANISE dataset [36] is a synthetic human-scene interaction dataset. It is mainly built upon two primary datasets: the AMASS dataset [19] and the ScanNet dataset [6]. HUMANISE combines these datasets by incorporating different global motions to meet contact and collision constraints. Since the original split includes duplicated human motions in the training and test set which is not suitable for motion prediction, we create a new split to ensure that motions in the test set are never observed during training. Moreover, we ensure that the scenes paired with the motions in the test set are categorized into seen scenes, which include a set of scenes observed during the training, and unseen scenes consisting of scenes never observed during training. Our final split comprises approximately 20k sub-sequences for training and 6k for testing, which includes 5k sub-sequences from seen scenes and 1k from unseen scenes. The videos run at 30 frame-per-second (FPS). Both our model and baseline models are trained to observe past 15 frames (0.5 sec.) and predict future 30 frames (1 sec.) because the average length of motion sequences in the dataset is 2.2 sec. which is much shorter than 3 sec.
@@
GIMO. The Gaze-Informed Human Motion dataset (GIMO) [40] is a real human-scene interaction dataset. The dataset captures human motion using an IMU-based motion capture system, and the 3D scene is scanned using a smartphone equipped with lidar sensors. Same as [36], GIMO also adopts SMPL-X [25] representation. It captures 11 subjects performing various actions e.g., open window, lie in bed, in 19 scenes, resulting in 127 long motion sequences with 192k frames. We follow the official training and test split, resulting in around 6k training sequences and 1k testing sequences. Same as HUMANISE [36], frame rate is 30Hz. Our model and baseline models predict future 30 frames with 15 past frames.